OpenStudy (rational):

factor \(x^4+x^2+25\)

OpenStudy (crashonce):

let x^2 = y y^2+y+25

OpenStudy (rational):

that doesnt factor nicely

OpenStudy (crashonce):

does it have to be nice lol

OpenStudy (crashonce):

oh unreal roots

OpenStudy (rational):

haha no, just want to factor it

OpenStudy (rational):

wolfram gives \[x^4+x^2+25=(x^2+3x+5)(x^2-3x+5)\]

OpenStudy (rational):

which looks nice hm

OpenStudy (crashonce):

find its apex. its 0,25

OpenStudy (rational):

who is apex

OpenStudy (crashonce):

the apex of the curve is at 0,25

OpenStudy (crashonce):

it has no intercepts so thats annoying, appart from the apex that is

OpenStudy (ikram002p):

I only can see trying lol x^4 and 25 are perfect square so we wanna find combination that eliminates other turms

OpenStudy (ikram002p):

(x^2+d1+5)(x^2+d2+5)

OpenStudy (rational):

thats how i would try but there seems to be some trick to factor this

OpenStudy (rational):

it looks special to me but idk if the special tricks are any useful because they don't work in general

OpenStudy (ikram002p):

Hmmm its like a ring

OpenStudy (ikram002p):

Lets try to factor this X^4+x^2+9 Seems doesn't factor nice same way Lets see X^4+x^2+64 Same

OpenStudy (ikram002p):

Hmmm so if we have x^4+x2+n^2 It can be factorized nicely as same ur question if 2n_1 is perfect square

OpenStudy (ikram002p):

I mean 2n minus 1

OpenStudy (ikram002p):

!_!

OpenStudy (sdfgsdfgs):

(x^2+d1+5)(x^2+d2+5) but since there is no x^3 term, d1=-d2 then looking at the x^2 term by expansion, it is 5 + 5 + d(-d) and equal to 1 so d^2=10-1 and d=3 In general (x^2+dx+n)(x^2-dx+n) will always expend to x^4 + (2n-d^2)*x^2 + n^2 Not so easy to spot but once u see it, u can recognize its symmetry.

OpenStudy (ikram002p):

Thats what I thought by saying a ring

OpenStudy (rational):

I think that should be the general method! this is a shortcut im talking about but there is no guarantee that this works always \[\begin{align}x^4+\color{blue}{x^2}+25&= x^4+\color{blue}{10x^2-9x^2}+25\\ &=x^4+10x^2+25 - 9x^2\\ &= (x^2+5)^2-(3x)^2\\ &=(x^2+5+3x)(x^2+5-3x) \end{align}\]

OpenStudy (ikram002p):

Nice !!!